Solve by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
2x + 3y = 1
5x + 3y = 16

Answer by checkley71(8403)

2X+3Y=1
5X+3Y=16 SUBTRACTING WE GET
------------
-3X=-15 OR X=15/3 OR X=5 & 2*5+3Y=1 OR 10+3Y=1 OR 3Y=-9 OR Y=-9/3 OR Y=-3